OpenStudy (anonymous):

how do you find the GCF of 24 and 32?

OpenStudy (anonymous):

factor both terms such that 24=2*2*2*3 = \[2^{3}*3\] 32=2*2*2*2*2 = \[2^{5}\] Then compare the like terms and pick the largest factor.

OpenStudy (anonymous):

that is, the largest factor common to both numbers.

OpenStudy (anonymous):

thanks for the help!